OpenStudy (anonymous):

solve for X 3Ln(2X+3)-2=1

OpenStudy (anonymous):

3ln (2x+3)-2 = 1 3 ln (2x+3) = 1+ 2 Ln (2x +3)=3/3 (2x+3)= e^1 X= (e-3)/2